Write an equation for the parabola with focus (1,3) and vertex (0,3).
Thanks

Answer by venugopalramana(3286)

Write an equation for the parabola with focus (1,3) and vertex (0,3).
Thanks
STD.EQN.IS
(Y-K)^2=4A(X-H)...WHERE
VERTEX =(0,3)=(H,K)......H=0,K=3
FOCUS =(1,3)=(H+A,K)......H+A=1...0+A=1....A=1
HENCE EQN. IS
(Y-3)^2=4X